OpenStudy (anonymous):

f(-7)=4(-7)^2+8(-7)-5 help! i have forgotten the basic algebra rules...

OpenStudy (anonymous):

f(-7)-135=0

OpenStudy (anonymous):

the answer should be 135, according to mathlab. i don't get that.

OpenStudy (lgbasallote):

4(-7)^2+8(-7)-5 4(49) - 56 - 5 196 - 56 - 5 191 - 56 =135

OpenStudy (anonymous):

dealing with functions.. f(z)=4z^2+8z-5, when input is -7

OpenStudy (mysesshou):

f(-7)=4(-7)^2+8(-7)-5 here, exponents happen first so, f(-7)=4*49+8(-7)-5 then multiplication (and division) f(-7)=196-56-5 eh, lgbasallote types faster than me. :)

OpenStudy (lgbasallote):

hahaha :))) good game ^_^

OpenStudy (anonymous):

just one question: how does 4(-7)^2 become positive 196?

OpenStudy (lgbasallote):

4(49)

OpenStudy (mysesshou):

exponents are done first in the order of operations. Then the multiplication by 4

OpenStudy (mysesshou):

Order: ( ) exponents * and / + and - etc. etc.

OpenStudy (anonymous):

yes, understood, so the neg squared becomes pos. My calculator confused me, as it shows -49.

OpenStudy (mysesshou):

try putting the ( ) around the -7 when you put it to the power.

OpenStudy (mysesshou):

so, type in ( -7 ) ^2

OpenStudy (anonymous):

Ok! Thx to both of you for your help:)

OpenStudy (mysesshou):

welcome :)